# An equivalence in random matrix and tensor models via a dually weighted intermediate field representation

### v1

@@ 3.3 Motivating the equivalence with a complex matrix model
With only one vertex type present, the generated graphs form a necklace-like sequence of loops of identical quartic vertices (See Fig. 2). These diagrams lack the $AAB$-vertices and face constraints necessary to realize the CDT structure. Thus, (3.10) does not reproduce CDT—but it illustrates how making a change of variables to move the external matrix $C_{2}$ into the propagator dramatically alters the combinatorial structure of the Feynmann expansion.
@@ 4 Equivalence in random matrix models
This section establishes an exact correspondence between two seemingly different matrix models: (i) a complex matrix model with covariance $(P,Q)$ and potential $V(M^{\dagger}M)$, and (ii) a self-adjoint two-matrix model whose kinetic term is the mixed bilinear ${\rm Tr}(AB)$ and whose $B$-field carries a logarithmic potential. The equivalence is exact for every choice of the potential $V$, and it holds at the level of full partition functions and all observables depending only on $M^{\dagger}M$ (or equivalently on $A$). In summary, this result shows that a complex matrix integral with general covariance can be reformulated as a self-adjoint two-matrix model whose intermediate field carries a dually weighted logarithmic interaction.
@@ 4.1 The complex matrix model
define a complex matrix model with arbitrary potential $V$, usually taken analytic in the self-adjoint matrix $M^{\dagger}M$. Expectation values in this model are defined by
@@
Thus a contraction ‘carries’ the matrix $Q$ along the first index and $P$ along the second one, giving rise to two complementary families of faces in the ribbon-graph expansion (See Figure 5).
@@
A schematic depiction of the two–stranded propagator and the two families of faces appears in Figure 4.
@@
$e^{V(A)}=\sum_{n=0}^{\infty}\;\sum_{\sigma\in S_{n}}\frac{1}{n!}\,w_{[\sigma]}{\rm Tr}_{[\sigma]}(A)\;.$ (4.5)
Substituting this expression into (4.1) yields
@@
Each term in the triple sum represents a ribbon graph with two families of faces weighted respectively by $Q$ and $P$ as illustrated in Figure 4.
@@ 4.2 The self-adjoint two–matrix model
Thus the only edges of the ribbon graphs appearing in the Feynman expansion are $AB$-propagators. This structure will be crucial since it is exactly this mixed pairing that allows $B$ to act as an intermediate field imposing constraints on $A$.
@@
A simple example is the Gaussian choice $Y[P,Q](B)=Y_{2}[P](B):=-\tfrac{1}{2}\,N\,{\rm Tr}((PB)^{2})$, for which (4.10) becomes
@@
In this case, the integral over $B$ can be performed explicitly, giving
@@
$Y[P,Q](B)=-{\rm Tr}\ln\left(\mathbb{1}^{\otimes 2}-i\,Q\otimes(PB)\right).$ (4.14)
Its Gaussian expectations reproduce exactly the two–face structure of the complex matrix model:
@@
Let $A\in{\rm H}(N)$ be distributed according to $\mathcal{Z}_{\rm HM}[0,Y[P,Q]]$ (4.7). For any $\sigma\in S_{n}$,
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y[P,Q]]=N^{-n}\sum_{\gamma\in S_{n}}{\rm Tr}_{[\gamma]}(Q)\,{\rm Tr}_{[\gamma\sigma]}(P).$ (4.15)
@@
It is straightforward to note that the right-hand side of (4.15) is identical to the expression obtained for the complex matrix model in Proposition 4.1. Thus, the two theories (4.1) and (4.7) assign the same weights to every multi-trace observable in the Gaussian sector. This strongly suggests a deeper correspondence that goes beyond expectation values. We show that this is not merely a formal similarity: the two partition functions are exactly equivalent. The matrix $B$ functions as the intermediate field whose logarithmic potential encodes the covariance of the complex matrix model (4.1), and the identification of both theories follows from integrating out the resulting Gaussian distribution of $M$ from (4.1).
@@ 4.3 Exact matrix equivalence via intermediate fields
The previous subsections 4.1 and 4.2 showed that for $V=0$, the complex and self-adjoint models match at the level of all multi-trace observables. We show that this agreement is an implication of a more general result: the two full partition functions ($V\neq 0$) coincide exactly. That is, the partition function for complex matrices (4.1) and the partition function for self-adjoint matrices (4.7) with logarithmic potential (4.14) coincide. This property has been observed in [difrancesco1992generatingfunctionfatgraphs] for a quadratic potential. Furthermore, the bridge between the two theories is provided by an intermediate–field representation, in which the self-adjoint matrix $B$ of (4.7) with (4.14) plays the role of a Fourier variable enforcing the constraint $A=M^{\dagger}M$. Once this constraint is made explicit, the logarithmic potential $Y[P,Q](B)$ (4.14) naturally emerges from integrating out $M$ from (4.1).
@@
Let $\mathcal{Z}_{\rm CM}[V](P,Q)$ (4.1) be the partition function for the complex matrix model, and $\mathcal{Z}_{\rm HM}[V,Y]$ (4.7) be the partition function for the self-adjoint two–matrix model. Then for $Y[P,Q](B)$ as given in (4.14) one has that
$\frac{\mathcal{Z}_{\rm CM}[V](P,Q)}{\mathcal{Z}_{\rm CM}[0](P,Q)}=\frac{\mathcal{Z}_{\rm HM}[V,Y[P,Q]]}{\mathcal{Z}_{\rm HM}[0,Y[P,Q]]}\;.$ (4.16)
@@
We note that the integral over $M$ is Gaussian which is clear in tensor notation,
@@
The remaining integral over $A$ and $B$ is precisely $\mathcal{Z}_{\rm HM}[V,Y[P,Q]]$, where the intermediate field $B$ has a dually weighted potential. The constant ${\cal N}_{\rm HM}^{-1}{\cal N}_{\rm CM}$ does not depend on $V$, thus the proof is complete. ∎
@@
Theorem 4.3 shows that, after integrating out the complex matrix $M$, the self-adjoint intermediate field, $B$, acquires a logarithmic potential of the dually weighted type. The intermediate field serves as a link between the two matrix models, transforming the two–index-cycle structure of the complex one-matrix model into a self-adjoint two-matrix model.
@@
In summary, in the complex matrix model $\mathcal{Z}_{\rm CM}$, the Gaussian measure couples $Q$ to the first index of $M$ and $P$ to the second. As a consequence, the ribbon graphs generated by the theory naturally carry two independent families of faces: one tracking the index cycle acted on by $Q$, and one tracking the index cycle acted on by $P$. These two families of faces coexist within a single connected Feynman graph and are the geometric origin of the weights appearing in Proposition 4.1. In the self-adjoint matrix model $\mathcal{Z}_{\rm HM}$, the interaction produced by the intermediate field $B$ reproduces exactly the same structure: the propagator determines one family of faces, while the dually weighted potential $Y[P,Q]$ supplies the weights for the other. By convention, in our presentation, the QQ–dependent contribution appears as a ribbon-graph vertex weight and the PP–dependent contribution as a ribbon-graph face weight, but this is a matter of choice: the roles of $P$ and $Q$ may be exchanged without changing the underlying combinatorics. Figures 3 and 4 illustrate this correspondence. Figure 3 shows how elementary ribbon-graph elements (edges, corners, and faces) are mapped between the two-matrix model formulations, while Figure 4 displays an example of the correspondence in a full ribbon graph, where the complex model is recovered from the self-adjoint theory with the dually weighted potential.
Figure 3: An example of equivalence between graph elements in the complex and the self-adjoint matrix models. On the rightmost columm, where a quadratic and a cubic vertices are shown, the red shapes represent a factor of ${\rm Tr}(Q^{2})$ and ${\rm Tr}(Q^{3})$, respectively.
Figure 4: Left: An example of a graph in the complex matrix theory (4.1) with particular potential $V$ including ${\rm Tr}((M^{\dagger}M)^{2})$ and ${\rm Tr}((M^{\dagger}M)^{3})$. A factor of $Q$ (resp. $P$) in the propagator is represented by blue (resp. red) lines, and interaction vertices ${\rm Tr}((M^{\dagger}M)^{2})$ and ${\rm Tr}((M^{\dagger}M)^{3})$ are represented in black. Right:An example of equivalence between a graph in the complex and the self-adjoint matrix models.
Expectation values
@@
$\left\langle f(M^{\dagger}M)\right\rangle[V](P,Q)=\left\langle f(A)\right\rangle[V,Y[P,Q]].$ (4.26)
@@
$\left\langle{\rm Tr}_{[\sigma]}(M^{\dagger}M)\right\rangle[0](P,Q)=\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y[P,Q]].$ (4.27)

### v2

@@ 3.3 Motivating the equivalence with a complex matrix model
With only one vertex type present, the generated graphs form a necklace-lik sequence of loops of identical quartic vertices (See Fig. 2). These diagrams lack the $AAB$-vertices and face constraints necessary to realize the CDT structure. Thus, (3.10) does not reproduce CDT.
@@
4 Summary of main results
In this section, we survey the main results of this paper.
Matrix models.
$\mathcal{Z}_{\rm HM}[V,{Y}_{\ln}]=\iint_{{\rm H}(N)^{2}}{\cal D}A\,{\cal D}B\;e^{-iN{\rm Tr}(AB)+{Y}_{\ln}[P,Q](B)+V(A)},\hskip 2.84526pt{Y}_{\ln}[P,Q](B)=-{\rm Tr}\ln\left(\mathbb{1}^{\otimes 2}-i\,Q\otimes(PB)\right).$ (4.2)
The intermediate effective field $A$ carries the effective potential $V(A)$, while the dually-weighted intermediate field $B$ carries a dually-weighted logarithmic potential $Y_{\ln}[P,Q](B)$.
Then, Theorem 5.3 implies that: for any potential $V$,
$\frac{\mathcal{Z}_{\rm CM}[V](P,Q)}{\mathcal{Z}_{\rm CM}[0](P,Q)}=\frac{\mathcal{Z}_{\rm HM}[V,{Y}_{\ln}[P,Q]]}{\mathcal{Z}_{\rm HM}[0,{Y}_{\ln}[P,Q]]}.$ (4.3)
In general, this equality holds at the level of formal power series and its convergence depends on the form of the potential VV [^7] The convergence criteria of $Z_{\rm CM}$ in terms of the spectra of $P$ and $Q$ is studied in Appendix C..
Equivalently, any observable depending only on the self–adjoint combination $M^{\dagger}M$ can be computed in either model. Corollary 5.3.1 states that, for any potential $V$ and any function $f:{\rm H}(N)\to\mathbb{C}$,
$\left\langle f(M^{\dagger}M)\right\rangle[V](P,Q)=\left\langle f(A)\right\rangle[V,{Y}_{\ln}[P,Q]].$ (4.4)
So far, the equivalence has been stated algebraically. A complementary combinatorial perspective on this equivalence follows from Propositions 5.1 and 5.2 (later contained in Corollary 5.3.1), which state that all multi-trace observables match in the two models. In the complex matrix model, multi-trace observables correspond to ribbon graphs with two complementary families of faces, weighted by the matrices $P$ and $Q$, respectively (Proposition 5.1). In contrast, Proposition 5.2 shows that in the self-adjoint formulation this structure is reorganized into a dually weighted model, in which faces carry weights associated with $P$, while vertices are weighted by $Q$. The equivalence can thus be viewed as a reorganization of the same combinatorial data, trading one set of face weights for vertex weights through the introduction of the intermediate field.
Tensor models.
The equivalence between complex and self-adjoint theories persists also in tensor models. Let $\varphi\in{\mathbb{C}}_{N}^{\otimes D}$ be a random complex order–$D$ tensor with covariance $R\in{\mathbb{C}}_{N}^{\otimes D}\otimes{\mathbb{C}}_{N}^{*\otimes D}$, and $V(\varphi\varphi^{\dagger})$, a scalar-$U(1)$ invariant potential, as elaborated in Subsection 6.1. As in (6.8), the partition function of the complex theory in question is
We consider random tensors $\Phi,\Psi\in{\rm H}({\mathbb{C}}_{N}^{\otimes D})$, both serving as intermediate fields, distributed according to the partition function of the self-adjoint theory (6.15) with potential (6.24)
${\cal Z}_{\rm HT}[V,{Y}_{\ln}[R]]=\iint{\cal D}\Phi{\cal D}\Psi\;e^{-i{\rm Tr}(\Phi\Psi)+{Y}_{\ln}[R](\Psi)+V(\Phi)}\,,\qquad{Y}_{\ln}[R](\Psi)=-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR\Psi\right)\,$, (4.6)
where the potential $Y[R](\Psi)$ gives $\Psi$ the role of the dually-weighted intermediate field. Our Theorem 6.3 states that, for any potential function $V$,
$\frac{\mathcal{Z}_{\rm CT}[V](R)}{\mathcal{Z}_{\rm CT}[0](R)}=\frac{\mathcal{Z}_{\rm HT}[V,Y_{\ln}[R]]}{\mathcal{Z}_{\rm HT}[0,Y_{\ln}[R]]}\;.$ (4.7)
Just as in the matrix model case, this should be understood as an equality between the formal power series, since its convergence depends on the potential $V$. Equivalently, as stated in Corollary 6.3.1, for any potential $V$ and any function $f:{\rm H}({\mathbb{C}}_{N}^{\otimes D})\to\mathbb{C}$,
$\left\langle f(\varphi\varphi^{\dagger})\right\rangle[V](R)=\left\langle f(\Phi)\right\rangle[V,{Y}_{\ln}[R]].$ (4.8)
Under an additional symmetry, i.e., the partial-$U(N)$ invariance explained in Subsection 6.1, the self–adjoint tensor model can be reduced to a model with tensors of order 2(D1)2(D-1) (6.40), as shown in Theorem 6.4. By working with the partial trace ${\rm Tr}_{(1)}(\varphi\varphi^{\dagger})$, this takes the form
$\frac{\mathcal{Z}_{\rm CT}[\hat{V}\circ{\rm Tr}_{(1)}](R)}{\mathcal{Z}_{\rm CT}[0](R)}=\frac{\mathcal{Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}_{\ln}[R]]}{\mathcal{Z}_{\rm\hat{H}T}[0,\hat{Y}_{\ln}[R]]}\;.$ (4.9)
We remark that when $D=2$ and the propagator factorizes as $R=Q\otimes P^{{\rm T}}$, the tensor equivalence of Theorem 6.4 has the matrix equivalence in Theorem 5.3 as a direct corollary.
In summary, our results presented in Theorems 5.3, 6.3, and 6.4 show that complex matrix and tensor models with nontrivial covariance and interactions depending on $M^{\dagger}M$ or $\varphi\otimes\varphi^{\dagger}$, respectively, admit exact self–adjoint representations in which the covariance data is carried by a dually-weighted intermediate field $B$ or $\Psi$, respectively.
This equivalence provides both a conceptual reinterpretation of such models and a practical computational tool: any expectation values/observable built from the self–adjoint combination may be computed in whichever formulation (complex or self–adjoint). In particular, self-adjoint models have lower order potential, so they tend to provide easier computation. In certain simple cases, the complex model in question may even correspond to a Gaussian self-adjoint models. In this work, we give an example of this simplification in the case of the CDT-like matrix model in (5.41) and in Appendix E.
Relation to existing intermediate field representations.
where ${\cal N}=(\mathcal{Z}_{\rm HT}[0,Y[R]])^{-1}(\mathcal{Z}_{\rm CT}[0](R))$, and, with an additional partial symmetry (Theorem 6.4),
for when there is a $\hat{V}$ such that $V(\varphi\varphi^{\dagger})=\hat{V}({\rm Tr}_{(1)}(\varphi\varphi^{\dagger}))$. Previously known intermediate field representations arise from suitable specializations of this identity (4.10).
    In [9], the intermediate field formula is developed for a single trace-invariant tensor interaction ${\rm Tr}_{B}(T,\bar{T})$ with unit covariance and is proven as a formal series in the coupling constant $\lambda$. In our notation, their intermediate field formula corresponds to choosing $R={\mathbb{1}}$ and repeatedly introducing a new field $\Psi_{i}$ for each pair of the $k$ pairs of $\varphi$ with $\varphi^{\dagger}$ in a 2k-vertex bubble $B$, so $V(\Phi)=\frac{\lambda}{n}N^{s}\,V_{B,\Omega}(\{\Phi_{i}\})$ (Eq. (52) in [9]). Making these substitutions in (4.10), we obtain the following representation (Eq. (53) in [9])
    $Z_{B}(\lambda,N)=\iint\prod_{i=1}^{k}{\cal D}\Phi_{i}\,{\cal D}\Psi_{i}\;e^{-i{\rm Tr}(\Psi\Phi)-{\rm Tr}\ln({\mathbb{1}}^{\otimes D}-i\Psi)+\frac{\lambda}{n}N^{s}V_{B,\Omega}(\Phi)}.$ (4.12)
    A transformation similar to a Wick rotation, turning the self-adjoint tensors $(\Psi,\Phi)$ into a complex adjoint pair $(M,M^{\dagger})$ yields the formulation that appears in [9]. However, a conceptual difference is that our equality holds exactly, while the derivation in [9] is performed at the level of formal power series.
  • Hubbard–Stratonovich representation.
    The classical intermediate field representation for the quartic one-matrix model, see e.g. [24], follows from choosing $D=2$, $R=\sqrt{\frac{\lambda}{N}}{\mathbb{1}^{\otimes 2}}$, and $\hat{V}(\hat{\Phi})=-\frac{1}{2}{\rm Tr}(\hat{\Phi}^{2})$. Our identity (4.11), after integrating out $\Phi$, yields
In summary, the various intermediate field constructions developed in the matrix and tensor literature arise as specific choices of the covariance $R$ and the potential $V$ in the universal identity (4.10) that we newly discovered in this paper. This provides a unified analytic framework that extends these previous results to completely general potentials and covariances.
@@ 5 Equivalence in random matrix models
This section establishes an exact correspondence between two seemingly different matrix models: (i) a complex matrix model with covariance $(P,Q)$ and potential $V(M^{\dagger}M)$, and (ii) a self-adjoint two-matrix model whose kinetic term is the mixed bilinear ${\rm Tr}(AB)$ and whose dually-weighted intermediate $B$-field carries a logarithmic potential. The equivalence between the partition functions of these models, as well as all observables depending only on $M^{\dagger}M$ (or equivalently on the intermediate effective field $A$), holds for arbitrary potentials $V$. In general, this equivalence is to be understood at the level of formal power series, while for suitable choices of $V$ it also holds at the level of convergent integrals. In summary, this result shows that a complex matrix integral with covariance of a general form can be reformulated as a self-adjoint two-matrix model whose intermediate field $B$ carries a dually weighted logarithmic interaction.
@@ 5.1 The complex matrix model
define a complex matrix model with arbitrary potential $V$, usually taken analytic in the self-adjoint matrix $M^{\dagger}M$. Expectation values in this model can be computed by
@@
Thus a contraction ‘carries’ the matrix $Q$ along the first index and $P$ along the second one, giving rise to two complementary families of faces in the ribbon-graph expansion.
Figure 3: Ribbon graph representation of the propagator of the complex matrix model defined by the partition function (5.1) and given in (5.3). The blue and red strands respectively represent the presence of $P$ and $Q$.
(a) ${\rm Tr}(M^{\dagger}M)$ vertex.
(b) ${\rm Tr}(M^{\dagger}M)^{2}$ vertex.
(c) ${\rm Tr}(M^{\dagger}M)^{3}$ vertex.
Figure 4: Examples of interaction vertices that can be considered in a theory of the form given in (5.1) via ribbon graph representation.
(a) Length-1 face.
(b) Length-2 face.
(c) Length-3 face.
Figure 5: Examples of faces that can be generated in (5.1) when vertices (see Fig. 4) are connected by propagators (see Fig. 3).
@@
A schematic depiction of the two–stranded propagator and the two families of faces appears in Figure 5 locally for the “QQ–faces” and Figure 10(a) in a given example of a ribbon graph.
@@
$e^{V(A)}=\sum_{n=0}^{\infty}\;\sum_{\sigma\in S_{n}}\frac{1}{n!}\,w_{[\sigma]}{\rm Tr}_{[\sigma]}(A)\;,$ (5.5)
where $w$ is a complex parameter for a given equivalence class of an symmetric group element. Substituting this expression into (5.1) yields
@@
Each term in the triple sum represents a ribbon graph with two families of faces weighted respectively by $Q$ and $P$ as illustrated in Figure 10(a).
@@ 5.2 The self-adjoint two–matrix model
Thus the only edges of the ribbon graphs appearing in the Feynman expansion are $AB$-propagators. This propagator is illustarted in Figure 6.
Figure 6: Ribbon graph representation of the propagator (5.9) of the self-adjoint matrix model defined by the partition function (5.15).
@@
A simple example is the Gaussian choice $Y(B)=Y_{2}[P](B):=-\tfrac{1}{2}\,N\,{\rm Tr}((PB)^{2})$, for which (5.10) becomes
@@
In this case, the integral over the field $B$ can be performed explicitly, giving
@@
(a) ${\rm Tr}\,A$ vertex.
(b) ${\rm Tr}\,A^{2}$ vertex.
(c) ${\rm Tr}\,A^{3}$ vertex.
Figure 7: Ribbon graph representation of various examples of interactions of the field $A$ in the partition function (5.15).
(a) ${\rm Tr}Q\;{\rm Tr}(PB)$ vertex.
(b) ${\rm Tr}Q^{2}\;{\rm Tr}(PB)^{2}$ vertex.
(c) ${\rm Tr}Q^{3}\;{\rm Tr}(PB)^{3}$ vertex.
Figure 8: Ribbon graph representation of various examples of interactions of the dually-weighted field $B$ in the partition function (5.15). The blue corners represent a factor of $P$ in $(PB)^{n}$, and the red loops represent a factor of ${\rm Tr}\,Q^{n}$.
@@
$Y_{\ln}[P,Q](B)=-{\rm Tr}\ln\left(\mathbb{1}^{\otimes 2}-i\,Q\otimes(PB)\right)=\sum_{k=1}^{\infty}\frac{i^{k}}{k}{\rm Tr}Q^{k}\,{\rm Tr}(PB)^{k},$ (5.14)
leading to the partition function
$\mathcal{Z}_{\rm HM}[V,Y_{\ln}[P,Q]]=\iint_{{\rm H}(N)^{2}}{\cal D}A\,{\cal D}B\;e^{-iN{\rm Tr}(AB)+V(A)-{\rm Tr}\ln(\mathbb{1}^{\otimes 2}-i\,Q\otimes(PB))}.$ (5.15)
Figures 7 and 8 illustrate several interactions that can be considered in this model. The Gaussian expectations reproduce exactly the two–face structure of the complex matrix model:
@@
Let $A\in{\rm H}(N)$ be distributed according to $\mathcal{Z}_{\rm HM}[0,Y_{\ln}[P,Q]]$ (5.7). For any $\sigma\in S_{n}$,
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y_{\ln}[P,Q]]=N^{-n}\sum_{\gamma\in S_{n}}{\rm Tr}_{[\gamma]}(Q)\,{\rm Tr}_{[\gamma\sigma]}(P).$ (5.16)
@@
It is straightforward to note that the right-hand side of (5.16) in Proposition 5.2 is identical to the expression obtained for the complex matrix model in Proposition 5.1. Thus, the two theories (5.1) and (5.7) assign the same weights to every multi-trace observable in the Gaussian sector ($V=0$). This strongly suggests a deeper correspondence that goes beyond expectation values. We show that this is not merely a formal similarity: the two partition functions are exactly equivalent. The matrix $B$ functions as the intermediate field whose logarithmic potential encodes the covariance of the complex matrix model (5.1), and the identification of both theories follows from integrating out the resuling Gaussian distribution of $M$ from (5.1).
@@ 5.3 Exact matrix equivalence via intermediate fields
The previous subsections 5.1 and 5.2 showed that for $V=0$, the complex and self-adjoint models match at the level of all multi-trace observables. We show that this agreement is an implication of a more general ($V\neq 0$) result: the partition function for complex matrices (5.1) and the partition function for self-adjoint matrices (5.7) with logarithmic potential (5.14) coincide. This property has been observed in [15] for a quadratic potential. Furthermore, the bridge between the two theories is provided by an intermediate–field representation, in which the self-adjoint matrix $B$ (i.e., the dually-weighted intermediate field) of (5.15) plays the role of a Fourier variable enforcing the constraint $A=M^{\dagger}M$. Once this constraint is made explicit, the logarithmic potential $Y_{\ln}[P,Q](B)$ (5.14) naturally emerges from integrating out $M$ from (5.1). We remark that, depending on the potentials $V$ and $Y$, the partition functions (5.1) and (5.15) may not be convergent. In that case, one may only be interested in their representation as a formal power series, or as a generating functions of discrete geometries.
@@
Let $\mathcal{Z}_{\rm CM}[V](P,Q)$ (5.1) be the partition function for the complex matrix model, and $\mathcal{Z}_{\rm HM}[V,Y]$ (5.7) be the partition function for the self-adjoint two–matrix model. Then, for $Y=Y_{\ln}[P,Q]$, as given in (5.14), one has that
$\frac{\mathcal{Z}_{\rm CM}[V](P,Q)}{\mathcal{Z}_{\rm CM}[0](P,Q)}=\frac{\mathcal{Z}_{\rm HM}[V,Y_{\ln}[P,Q]]}{\mathcal{Z}_{\rm HM}[0,Y_{\ln}[P,Q]]}\;$ (5.17)
as formal power series.
@@
For sufficiently well-behaved potentials $V$, for which the integrals are absolutely convergent, one may exchange the order of integration and perform the integral over $M$ first. More generally, for arbitrary potential $V$, the same manipulation can be carried out at the level of formal power series. In either case, the integral over $M$ is Gaussian, as is most clearly seen in tensor notation,
@@
The remaining integral over $A$ and $B$ is precisely $\mathcal{Z}_{\rm HM}[V,Y_{\ln}[P,Q]]$, where the intermediate field $B$ has a dually weighted potential. The constant ${\cal N}_{\rm HM}^{-1}{\cal N}_{\rm CM}$ does not depend on $V$, thus the proof is complete. ∎
@@
Theorem 5.3 shows that, after integrating out the complex matrix $M$, the self-adjoint intermediate field $B$ acquires a logarithmic potential of the dually weighted type. The intermediate field thus provides a link between the two matrix models, transforming the two–index-cycle structure of the complex one-matrix model into a self-adjoint two-matrix model.
Remark. The result is obtained at the level of formal power series since, depending on the choice of potential $V$, the partition functions involved may not be absolutely convergent or may only be defined in a distributional sense. In such situations, the equivalence is to be interpreted as an identity between formal expansions. However, for suitable choices of $V$ for which the integrals are well-defined (e.g. when the action is sufficiently bounded so that the integrals are absolutely convergent, or admit a well-defined Fourier transform), the equivalence holds at the level of convergent integrals. In this case, the derivation can be justified by exchanging the order of integration and performing the Gaussian integral over $M$ explicitly.
@@
In summary, in the complex matrix model $\mathcal{Z}_{\rm CM}$, the Gaussian measure couples $Q$ to the first index of $M$ and $P$ to the second. As a consequence, the ribbon graphs generated by the theory naturally carry two independent families of faces: one tracking the index cycle acted on by $Q$, and one tracking the index cycle acted on by $P$. These two families of faces coexist within a single connected Feynman graph and are the geometric origin of the weights appearing in Proposition 5.1. In the self-adjoint matrix model $\mathcal{Z}_{\rm HM}$, the interaction produced by the intermediate field $B$ reproduces exactly the same structure: the propagator determines one family of faces, while the dually weighted potential $Y_{\ln}[P,Q]$ supplies the weights for the other. By convention, in our presentation, the QQ–dependent contribution appears as a ribbon-graph vertex weight and the PP–dependent contribution as a ribbon-graph face weight, but this is a matter of choice: the roles of $P$ and $Q$ may be exchanged without changing the underlying combinatorics. We can visualize how elementary ribbon-graph elements (edges, corners, and faces) are mapped between the two-matrix model formulations with aid of Figures 3,  4,  5,  6,  7,  8 and  9. Vertices of $M$ (Figure 4) are equivalent to vertices of the intermediate effective field $A$ (Figure 7), and faces of $M$ (Figure 5) are equivalent to vertices of the dually-weighted intermediate field $B$ (Figure 8). Figure 10 illustrates this correspondence by displaying an example of the correspondence in a full ribbon graph, where the complex model is recovered from the self-adjoint theory with the dually weighted potential.
Figure 9: The interaction vertex ${\rm Tr}(BM^{\dagger}M)$, which appears in (5.21).
Figure 10: An example of equivalence between a graph in the: (a) complex matrix theory ${\cal Z}_{\rm CM}[V](P,Q)$ (5.1) using the graphical representations of the propagator, interaction vertices and faces shown in Figures 34, and 5; (b) intermediate theory (5.21) involving the propagators and interaction vertices shown in Figures 3, 5, 6, 7, and 9; (c) the Hermitian matrix theory $\mathcal{Z}_{\rm HM}[V,Y_{\ln}[P,Q]]$ (5.15) now composed of the propagator and interaction vertices shown in Figure 6, 7, and 8. Gray lines represent intermediate fields AA and $B$.
Expectation values.
@@
$\left\langle f(M^{\dagger}M)\right\rangle[V](P,Q)=\left\langle f(A)\right\rangle[V,Y_{\ln}[P,Q]].$ (5.27)
@@
$\left\langle{\rm Tr}_{[\sigma]}(M^{\dagger}M)\right\rangle[0](P,Q)=\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y_{\ln}[P,Q]].$ (5.28)